OpenStudy (anonymous):

x^4+2ix-3x^2-yi=3-5i+1+2iy what are the values of x and y

terenzreignz (terenzreignz):

First, simplify the right part. Might as well do the left part, too...

OpenStudy (anonymous):

x^4-3x^2=4 and 2x-3y+5=0

OpenStudy (anonymous):

plz simplify furthur

terenzreignz (terenzreignz):

You can solve for x in this equation \[\huge x^4-3x^2 = 4\]

terenzreignz (terenzreignz):

For which there are many possible values, possibly... \[\huge x^4 - 3x^2 - 4 = 0\]

terenzreignz (terenzreignz):

Oh, and before you panic, maybe you can let \[\huge u = x^2\]and solve for u instead...

terenzreignz (terenzreignz):

\[\huge u^2-3u-4 = 0\]

terenzreignz (terenzreignz):

@leena1996 work with me here.

terenzreignz (terenzreignz):

I do hope x is only allowed to take real number values, though....

OpenStudy (anonymous):

ya i got x=+/-2 and y should be 1/3 but i am getting it as 3....

terenzreignz (terenzreignz):

I wonder why... \[\huge 2x -3y + 5 = 0 \]\[\huge2(\pm2)-3y +5 = 0\]\[\huge \pm4-3y=-5\]\[\huge -3y=-5\mp4\] \[\huge y = \frac{5\pm4}{3}\]

terenzreignz (terenzreignz):

It seems both 1/3 and 3 are possible, @leena1996 depending on which value of x you pick ... If x = 2, then y = 3 If x = -2, then y = 1/3